Rewrite the expression as an algebraic expression in x.
tan(sin^-1(x))

Found 2 solutions by stanbon, ikleyn:
Answer by stanbon(75887)

write the expression as an algebraic expression in x.
tan(sin^-1(x))
-----
Tangent of the angle whose sin = x
------
By definition sin = opp/hyp; so opp = x and hyp = 1.
The adj = sqrt[x^2 + 1]
-------
Ans: Tan = opp/adj = x/sqrt[x^2+1]
---------------------
Cheers,
Stan H.
-----------

Answer by ikleyn(53751)

.
write the expression as an algebraic expression in x.
tan(sin^-1(x))
-----
Tangent of the angle whose sin = x
------
By definition sin = opp/hyp; so opp = x and hyp = 1.
The adj = sqrt(1-x^2)
-------
Ans: Tan = opp/adj = x/sqrt(1-x^2)